# Posterior Ramifications of Prior Dependence Structures

### v2

@@ 3 Objectives for Dependence Structure Specification
This section outlines several objectives that practitioners may wish to achieve when specifying prior distributions and their dependence structures. For each objective, we discuss whether retention of the prior dependence structure as defined in Corollary 1 is important. This knowledge can be used to assess the adequacy of an elicited prior distribution.
@@ 3.1 Supplementation of Small Samples
For this model with $\boldsymbol{\theta}=(\mu,\sigma^{2})$, the inverse Fisher information matrix $\mathcal{I}(\boldsymbol{\theta})^{-1}$ is diagonal for all possible values of $\mu\in\mathbb{R}$ and $\sigma^{2}\in\mathbb{R}^{+}$. It follows by Corollary 1 that the joint posterior of $\mu$ and $\sigma^{2}$ cannot retain a prior dependence structure for which Kendall’s $\tau$ is not 0. Under the conditions for Corollary 1, any attempt to inject positive or negative dependence into this posterior will be unsuccessful in the limit of infinite data. That is, once both $\mu$ and $\sigma^{2}$ are precisely estimated from the data, there will be no practical correlation between their estimates. However, imposing prior dependence is nevertheless useful for small sample sizes.
While chronically rejected dependence structures can supplement information from small samples, it would be ideal if the source of posterior dependence for $\boldsymbol{\theta}$ were transparent. Corollary 1 can be used to determine whether the prior or likelihood function gives rise to this dependence. This knowledge would provide additional context with which to interpret posterior analyses. This example also suggests that it is not worthwhile to spend time eliciting complicated prior dependence structures that cannot be retained when collecting large samples.
The use of informative priors to improve model identifiability is related to supplementing small samples with prior information. Although unidentifiability arises in other contexts, the sample size is often small with respect to the complexity of models that cannot be fully identified. Previous contributions detailed how using informative priors to enhance model identifiability impacts model interpretation and the convergence of Markov chain Monte Carlo methods (Gelfand and Sahu, 1999; Eberly and Carlin, 2000; Gustafson, 2005). These contributions mainly considered informative marginal priors, and discussion of informative prior dependence structures was limited.
@@ 3.2 Coverage of Credible Sets
Prior dependence structures may also be specified to calibrate Bayesian credible sets. In this context, Gustafson (2012) referred to the data generating prior as “nature’s” prior. The data generation prior in our framework is $p_{D}(\boldsymbol{\theta})$ from (5). If the analysis prior $p(\boldsymbol{\theta})$ coincides with “nature’s” prior and the model $m(\hskip 1.0pt\cdot\hskip 1.0pt|\hskip 1.0pt\boldsymbol{\theta})$ is correctly specified, credible sets attain their nominal coverage (Gustafson, 2012). We let $\mathcal{J}_{\boldsymbol{\theta},1-\alpha}(\boldsymbol{y}^{{}_{(n)}})$ denote a credible set for $\boldsymbol{\theta}$ given the observed data $\boldsymbol{y}^{{}_{(n)}}$ with coverage $1-\alpha$. A popular choice for the credible set is that of highest posterior density (HPD). Credible sets attain their nominal coverage when
@@
We now investigate how the prior dependence structure impacts the calibration of credible sets for the multinomial example from Section 2.3. We used the same prior predictive distribution of $\boldsymbol{Y}^{{}_{(n)}}$ for this numerical study, joining the $\text{BETA}(20,40)$ prior for $Z_{1}$ and $\text{BETA}(30,30)$ prior for $Z_{2}$ with a Gaussian copula parameterized with Pearson’s $\rho=-0.9$. For each of 10000 simulation repetitions, we approximated the posterior of $\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}}$ as described in Section 2.3 with “nature’s” analysis prior. For each posterior, we approximated its 95% HPD set for $Z_{1}$ and $Z_{2}$ using two-dimensional kernel density estimation (Ripley, 2002). Empirical coverage was estimated as the proportion of simulation repetitions for which the parameter value $\boldsymbol{\theta}_{0}=(Z_{1,0},Z_{2,0})$ used to generate the multinomial data was contained in this HPD set. We implemented this process for $n=\{10^{1},10^{2},10^{3},10^{4},10^{5}\}$. We then repeated this process for analysis priors $p(\boldsymbol{\theta})$ that joined the marginal beta priors from “nature’s” prior with a Gaussian copula parameterized by Pearson’s $\rho=\{-0.95,-0.85,-0.80,\dots,0.95\}$. The results from this numerical study are visualized in Figure 3.
@@
As the sample size $n$ increases, the impact of prior dependence is reduced and the empirical coverage approaches the nominal coverage for all $\rho$ values considered. This conclusion is to be expected given the asymptotic theory discussed in Section 2. The insights drawn from Figure 3 are not solely applicable to chronically rejected prior dependence structures. We observed similar results for a numerical study with the gamma model involving analysis priors that do not satisfy the conditions for chronic rejection in Corollary 1. These results are detailed in Appendix B of the online supplement.
Given the results from this subsection, we suggest erring on the side of understating the strength of dependence structures a priori. This recommendation is not made to encourage practitioners to engineer credible sets with coverages that exceed their nominal values. This excess coverage may not be desirable and cannot persist as the sample size increases. We instead make this recommendation to mitigate the potential ramifications of overstating the strength of dependence structures on the calibration of credible sets. The notion of harmful priors has been discussed elsewhere (see e.g., Reimherr et al. (2021)). However, priors that are harmful to the calibration of credible sets are not necessarily harmful priors in alternative contexts.
@@ 3.3 Inference Regarding Dependence Structures
One may also conduct posterior analyses to draw conclusions about the dependence structure of $\boldsymbol{\theta}$. The dependence structure is also relevant when considering the posterior of a function of several parameters in $\boldsymbol{\theta}$. To achieve these objectives with posterior analysis, we recommend using Corollary 1 to determine whether the prior dependence structure can be retained. Per Table 1, we can draw vastly different conclusions about chronically rejected dependence structures given small and large samples obtained from the same data generation process. Corollary 1 provides context to help gauge the reliability of our conclusions for small samples. For large samples, Corollary 1 dictates if time should be invested eliciting a dependence structure based on whether it can be retained a posteriori.
Corollary 1 could also be used to determine whether it is sensible to make inferences about dependence structures. While copula-based priors may be able to accommodate flexible dependence structures, the statistical models chosen for the likelihood function may not have this capability. For the multinomial model, there is a single component for each observation of $\boldsymbol{y}^{{}_{(n)}}=\{y_{i}\}_{i=1}^{n}$. It would be difficult to specify a likelihood function that could accommodate complex dependence structures between the conditional multinomial probabilities given the available data. As such, Corollary 1 could diagnose philosophical issues with objectives for posterior analyses.
@@
where the copula density function is parameterized by $\boldsymbol{\upsilon}$ and $\boldsymbol{\theta}=(\lambda,\kappa,\boldsymbol{\upsilon})$. If $c(u_{1},u_{2};\boldsymbol{\upsilon})$ corresponds to the independence copula, the posterior correlation between $\lambda$ and $\kappa$ based on the inverse Fisher information will approach 0 as $n\rightarrow\infty$. More flexible dependence structures for $\lambda$ and $\kappa$ could be accommodated a posteriori given different choices for $c(u_{1},u_{2};\boldsymbol{\upsilon})$. In those settings, eliciting dependence structures for $\lambda$ and $\kappa$ could be worthwhile – even for large samples since the dependence structure might be retained.
@@ 3.5 Posterior Concentration
Recent work has suggested that the choice of prior dependence structure can expedite the convergence of the posterior around a fixed parameter value $\boldsymbol{\theta}_{0}$. In particular, Michimae and Emura (2022) suggested that joint priors based on Archimedean copulas lead to more accurate and concentrated posterior distributions in the context of Bayesian ridge regression. The accuracy and concentration of the posterior of $\boldsymbol{\theta}$ were considered via the total mean absolute error between the posterior median and a fixed parameter value $\boldsymbol{\theta}_{0}$. Michimae and Emura’s (2022) numerical studies showed that the posterior was more concentrated around $\boldsymbol{\theta}_{0}$ when the marginal priors for the regression coefficients were joined using (Archimedean) Clayton and Gumbel copulas (Nelsen, 2006) instead of more standard Gaussian copulas.
The asymptotic theory presented in Section 2 informally suggests the choice of prior dependence structure cannot give rise to increased posterior concentration for large samples. For small samples, further investigation into how the prior dependence structure prompts increased posterior concentration is required. This investigation would disclose whether improving posterior concentration is a sensible objective for prior dependence specification. These recent recommendations serve as one source of motivation to study the impact of the prior copula on the posterior distribution, which we do in Section 4.
@@ 4.1 Convergence of the Posterior Mode
For an arbitrary model, it may be challenging to correctly specify the dependence structure a priori. As such, we consider how the choice of copula for the prior distribution impacts the posterior. We suppose that two potential priors for $\boldsymbol{\theta}$, denoted $p_{1}(\boldsymbol{\theta})$ and $p_{2}(\boldsymbol{\theta})$, are defined as in (4) using the same marginal distributions $F_{1},\dots,F_{d}$ but different copula density functions $c_{1}(\boldsymbol{u})$ and $c_{2}(\boldsymbol{u})$. We require that both copula density functions are absolutely continuous and twice differentiable with respect to $\boldsymbol{u}=(u_{1},\dots,u_{d})$. To ensure the domain of the parameter space is not inadvertently restricted, the priors should be chosen such that $p_{1}(\boldsymbol{\theta})>0$ and $p_{2}(\boldsymbol{\theta})>0$ for all $\boldsymbol{\theta}\in\boldsymbol{\Theta}$.
We define posteriors $p_{1}(\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}})$ and $p_{2}(\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}})$ by combining the likelihood $L(\boldsymbol{\theta};\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}})$ for the model $m(\boldsymbol{y}^{{}_{(n)}}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{\theta})$ with $p_{1}(\boldsymbol{\theta})$ and $p_{2}(\boldsymbol{\theta})$, respectively. We summarize each posterior via its posterior mode for $\boldsymbol{\theta}$, denoted by $\tilde{\boldsymbol{\theta}}^{{}_{(k)}}=\text{arg\,max}_{{\boldsymbol{\theta}}}~{}p_{k}(\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}})$ for $k=1,2$. In this section, we consider the convergence of the posterior mode to a fixed value $\boldsymbol{\theta}_{0}$. For a given sample $\boldsymbol{y}^{{}_{(n)}}$, we compare the Euclidean distance between $\boldsymbol{\theta}_{0}$ and each posterior mode, denoted by $\mathcal{D}_{k}=\lVert\tilde{\boldsymbol{\theta}}^{{}_{(k)}}-\boldsymbol{\theta}_{0}\rVert_{2}$ for $k=1,2$. Theorem 2 indicates that we generally do not expect the choice of prior copula to impact whether the posterior mode is closer to $\boldsymbol{\theta}_{0}$ for large sample sizes $n$.
@@
$\begin{split}\frac{\partial}{\partial\theta_{j}}\bigg{[}\text{log}(c_{2}(\boldsymbol{u}))-\text{log}(c_{1}(\boldsymbol{u}))\bigg{]}=f_{j}(\theta_{j})\frac{\partial}{\partial u_{j}}\bigg{[}\text{log}(c_{2}(\boldsymbol{u}))-\text{log}(c_{1}(\boldsymbol{u}))\bigg{]}.\end{split}$ (8)
We can factor out the $f_{j}(\theta_{j})$ term because $p_{1}(\boldsymbol{\theta})$ and $p_{2}(\boldsymbol{\theta})$ are defined using the same marginals. The priors were also defined such that $p_{1}(\boldsymbol{\theta})>0$ and $p_{2}(\boldsymbol{\theta})>0$ for all $\boldsymbol{\theta}\in\boldsymbol{\Theta}$. Therefore, $f_{j}(\theta_{j})$ must be positive, and the partial derivative in (8) with respect to $\theta_{j}$ is 0 if and only if the partial derivative with respect to $u_{j}$ is 0. This correspondence ensures the results from Theorem 2 generalize over different specifications for the marginal distributions.

### v3

@@ 3 Objectives for Dependence Structure Specification
3.1 Overview
This section outlines several objectives that practitioners may wish to achieve when specifying prior distributions and their dependence structures. Our discussion is motivated by the following questions. When is it useful to encode prior dependence that is at odds with the asymptotic covariance structure? When is it sensible to draw inferences about the posterior dependence structure of $\boldsymbol{\theta}$? How are posterior inferences impacted by the dependence structure? Does the inability to retain prior dependence reveal philosophical issues with the objectives for posterior analysis?
The relevance of each question depends on the objective for prior dependence specification, but we consider retention of the prior dependence structure as defined in Corollary 1 for all objectives. The objectives explored in this section are not necessarily mutually exclusive. Nevertheless, we discuss why it may not be possible for a given prior dependence structure to facilitate certain combinations of objectives for posterior analysis. This discussion can guide conversations between statisticians and practitioners to simplify the prior specification process where possible and promote alignment between the prior dependence structure and posterior objectives.
@@ 3.2 Supplementation of Small Samples
For this model, the inverse Fisher information matrix $\mathcal{I}(\boldsymbol{\theta}_{0})^{-1}$ is diagonal for all possible values of $\boldsymbol{\theta}_{0}=(\mu_{0},\sigma^{2}_{0})=\mathbb{R}\times\mathbb{R}^{+}$. That is, $\tau(\boldsymbol{\theta}_{0})_{\mu,\sigma^{2}}=0$ for all $\boldsymbol{\theta}_{0}\in\boldsymbol{\Theta}*$. It follows by Corollary 1 that the joint posterior of $\mu$ and $\sigma^{2}$ cannot retain a prior dependence structure for which Kendall’s $\tau$ is not 0. Under the conditions for Corollary 1, any attempt to inject positive or negative dependence into this posterior will be unsuccessful in the limit of infinite data. Once both $\mu$ and $\sigma^{2}$ are precisely estimated from the data, there will be no practical correlation between their estimates. However, imposing prior dependence is nevertheless useful for small sample sizes.
When chronically rejected dependence structures are used to supplement information from small samples, we argue that drawing inferences about posterior dependence is problematic. These inferences would just confirm our beliefs about the prior dependence structure despite its incompatibility with likelihood function – and therefore the data. Such priors are specified for objectives that supersede learning about the dependence structure. For this example, the prior dependence structure was selected with the hope of obtaining a better marginal posterior for $\sigma^{2}$ with limited information. This example also suggests that it is not worthwhile to spend time eliciting complicated prior dependence structures that cannot be retained when collecting large samples. Discarding chronically rejected dependence structures can simplify the prior specification process for practitioners in large-sample settings.
The impact of posterior dependence on inference is of course not limited to conclusions about the dependence structure of $\boldsymbol{\theta}$. For instance, the dependence structure impacts the posterior of a function of the parameters in $\boldsymbol{\theta}$ and posterior predictive inference. In this example, the posterior of the 0.9 quantile of the $\mathcal{N}(\mu,\sigma^{2})$ distribution would be a function of both components in $\boldsymbol{\theta}$. This posterior incorporates information from the posterior dependence structure, but it also critically relies on the marginal posteriors of $\boldsymbol{\theta}$ that we prioritized over the dependence structure via prior specification. As such, we believe that these inferences are more philosophically consistent with our objectives for prior specification than those based exclusively on the dependence structure of $\boldsymbol{\theta}$.
@@ 3.3 Coverage of Credible Sets
Prior dependence structures may also be specified to calibrate Bayesian credible sets. In this context, Gustafson (Gustafson, 2012) referred to the data generating prior as “nature’s” prior. The data generation prior in our framework is $p_{D}(\boldsymbol{\theta})$ from (5). If the analysis prior $p(\boldsymbol{\theta})$ coincides with “nature’s” prior and the model $m(\hskip 1.0pt\cdot\hskip 1.0pt|\hskip 1.0pt\boldsymbol{\theta})$ is correctly specified, credible sets attain their nominal coverage (Gustafson, 2012). We let $\mathcal{J}_{\boldsymbol{\theta},1-\alpha}(\boldsymbol{y}^{{}_{(n)}})$ denote a credible set for $\boldsymbol{\theta}$ given the observed data $\boldsymbol{y}^{{}_{(n)}}$ with posterior probability of $1-\alpha$. A popular choice for the credible set is that of highest posterior density (HPD). Credible sets attain their nominal coverage when
@@
We now investigate how the prior dependence structure impacts inference via the calibration of credible sets for the multinomial example from Section 2.3. We used the same prior predictive distribution of $\boldsymbol{Y}^{{}_{(n)}}$ for this numerical study, joining the $\text{BETA}(20,40)$ prior for $Z_{1}$ and $\text{BETA}(30,30)$ prior for $Z_{2}$ with a Gaussian copula parameterized with Pearson’s $\rho=-0.9$. For each of 10000 simulation repetitions, we approximated the posterior of $\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}}$ as described in Section 2.3 with “nature’s” analysis prior. For each posterior, we approximated its 95% HPD set for $Z_{1}$ and $Z_{2}$ using two-dimensional kernel density estimation (Ripley, 2002). Empirical coverage was estimated as the proportion of simulation repetitions for which the parameter value $\boldsymbol{\theta}_{0}=(Z_{1,0},Z_{2,0})$ used to generate the multinomial data was contained in this HPD set. We implemented this process for $n=\{10^{1},10^{2},10^{3},10^{4},10^{5}\}$. We then repeated this process for analysis priors $p(\boldsymbol{\theta})$ that joined the marginal beta priors from “nature’s” prior with a Gaussian copula parameterized by Pearson’s $\rho=\{-0.95,-0.85,-0.80,\dots,0.95\}$. The results from this numerical study are visualized in Figure 3.
@@
As the sample size $n$ increases, the impact of prior dependence is reduced and the empirical coverage approaches the nominal coverage for all $\rho$ values considered. The insights drawn from Figure 3 are not solely applicable to chronically rejected prior dependence structures. We observed similar results for a numerical study with the gamma model involving analysis priors that do not satisfy the conditions for chronic rejection in Corollary 1. These results are detailed in Appendix B of the online supplement.
If a chronically rejected prior dependence structure is used with the intent of calibrating credible sets, we do not advise drawing inferences based solely on the posterior dependence structure. Prior dependence has been specified to prioritize the objective of calibration over the goal of learning about the dependence structure. Moreover, we suggest erring on the side of understating the strength of dependence structures a priori. This recommendation is not made to encourage practitioners to engineer credible sets with coverages that exceed their nominal values. This excess coverage may not be desirable and cannot persist as the sample size increases. We instead make this recommendation to mitigate the potential ramifications of overstating the strength of dependence structures on the calibration of credible sets. The notion of harmful priors has been discussed elsewhere (see e.g., (Reimherr et al., 2021)), but priors that are harmful to the calibration of credible sets are not necessarily harmful priors in alternative contexts.
@@ 3.4 Inference Regarding Dependence Structures
We have argued that drawing inference regarding posterior dependence is problematic when chronically rejected prior dependence structures are specified to achieve a superseding objective. However, drawing inference about the dependence structure may be our main posterior objective or equally as important as other objectives. In that case, we recommend using Corollary 1 to determine whether the prior dependence structure can be retained. It would be ideal if Bayesian inference informed by small and large samples could feasibly support our prior notions about the dependence structure. If the data via the likelihood function cannot possibly support these notions, we should ask practitioners whether the likelihood function or prior distribution require modification.
Corollary 1 could therefore help diagnose philosophical issues with posterior objectives. While copula-based priors may be able to accommodate flexible dependence structures, the statistical models chosen for the likelihood function may not have this capability. For the multinomial model, there is a single component for each observation of $\boldsymbol{y}^{{}_{(n)}}=\{y_{i}\}_{i=1}^{n}$. It would be difficult to specify a likelihood function that could accommodate complex dependence structures between the conditional multinomial probabilities given the available data. Thus, it is not always sensible to make inferences about dependence structures.
This lack of flexibility alludes to the trade-offs of using simple likelihoods instead of more realistic ones. For instance, Wilson (Wilson, 2018) considered the capacity of an engineering structure. These outcomes are influenced by several complex processes that include structural aging and the procedures to assess structure capacity. By representing these outcomes as multinomial, the complexity of the data generating process is simplified, which limits the scope of the conclusions that can be drawn. Gelman et al. (Gelman et al., 2017) argued that the prior can only be interpreted in the context of the likelihood function with which it is paired. This example also underscores the benefits of developing the prior and likelihood function in tandem.
@@
where the copula density function is parameterized by $\boldsymbol{\upsilon}$ and $\boldsymbol{\theta}=(\lambda,\kappa,\boldsymbol{\upsilon})$. If $c(u_{1},u_{2};\boldsymbol{\upsilon})$ corresponds to the independence copula, the posterior correlation between $\lambda$ and $\kappa$ based on the inverse Fisher information will approach 0 as $n\rightarrow\infty$. More flexible dependence structures for $\lambda$ and $\kappa$ could be accommodated a posteriori given different choices for $c(u_{1},u_{2};\boldsymbol{\upsilon})$. In those settings, eliciting dependence structures for $\lambda$ and $\kappa$ to support inference about posterior dependence could be worthwhile – even for large samples since the dependence structure might be retained.
@@ 3.6 Posterior Concentration
Recent work has suggested that the choice of prior dependence structure can expedite the convergence of the posterior around a fixed parameter value $\boldsymbol{\theta}_{0}$. This objective of increasing posterior concentration is related to the goal of improving model identifiability. The use of informative marginal priors to improve identifiability in complex settings has previously been discussed (Gelfand and Sahu, 1999; Eberly and Carlin, 2000; Gustafson, 2005).
In terms of the dependence structure, Michimae and Emura (Michimae and Emura, 2022) recently suggested that joint priors with vine structures based on Archimedean copulas lead to more accurate and concentrated posterior distributions in the context of Bayesian ridge regression. The accuracy and concentration of the posterior of $\boldsymbol{\theta}$ were considered via the total mean absolute error between the posterior median and a fixed parameter value $\boldsymbol{\theta}_{0}$. Michimae and Emura’s (Michimae and Emura, 2022) numerical studies showed that the posterior was more concentrated around $\boldsymbol{\theta}_{0}$ when the marginal priors for the regression coefficients were joined using (Archimedean) Clayton and Gumbel copulas (Nelsen, 2006) instead of more standard Gaussian copulas.
Under the conditions for the Bernstein-von Mises theorem, the asymptotic theory from Section 2 suggests the choice of prior dependence structure cannot give rise to increased posterior concentration for large samples. For small samples, however, further investigation into how the prior dependence structure prompts increased posterior concentration is required. This investigation would disclose whether improving posterior concentration is a sensible objective for prior dependence specification. As such, we study the prior copula’s impact on the posterior distribution in Section 4. This investigation considers the connection between posterior concentration and model identifiability and contextualizes our results within the broader discussion on the use of chronically rejected dependence structures.
@@ 4.1 Convergence of the Posterior Mode
Since it may be challenging to correctly specify the dependence structure a priori, we consider how the choice of copula for the prior distribution impacts the posterior. We suppose that two potential priors for $\boldsymbol{\theta}$, denoted $p_{1}(\boldsymbol{\theta})$ and $p_{2}(\boldsymbol{\theta})$, are defined as in (4) using the same marginal distributions $F_{1},\dots,F_{d}$ but different copula density functions $c_{1}(\boldsymbol{u})$ and $c_{2}(\boldsymbol{u})$. We require that both copula density functions are absolutely continuous and twice differentiable with respect to $\boldsymbol{u}=(u_{1},\dots,u_{d})$. To ensure the domain of the parameter space is not inadvertently restricted, the priors should be chosen such that $p_{1}(\boldsymbol{\theta})>0$ and $p_{2}(\boldsymbol{\theta})>0$ for all $\boldsymbol{\theta}\in\boldsymbol{\Theta}$.
We define posteriors $p_{1}(\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}})$ and $p_{2}(\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}})$ by combining the likelihood $L(\boldsymbol{\theta};\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}})$ for the model $m(\boldsymbol{y}^{{}_{(n)}}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{\theta})$ with $p_{1}(\boldsymbol{\theta})$ and $p_{2}(\boldsymbol{\theta})$, respectively. We summarize each posterior via its posterior mode for $\boldsymbol{\theta}$, denoted by $\tilde{\boldsymbol{\theta}}^{{}_{(k)}}=\text{arg\,max}_{{\boldsymbol{\theta}}}~{}p_{k}(\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}})$ for $k=1,2$. In this section, we consider the convergence of the posterior mode to a fixed value $\boldsymbol{\theta}_{0}$ because the posterior generally concentrates around its mode as data are observed. For a given sample $\boldsymbol{y}^{{}_{(n)}}$, we compare the Euclidean distance between $\boldsymbol{\theta}_{0}$ and each posterior mode, denoted by $\mathcal{D}_{k}=\lVert\tilde{\boldsymbol{\theta}}^{{}_{(k)}}-\boldsymbol{\theta}_{0}\rVert_{2}$ for $k=1,2$. Theorem 2 indicates that we generally do not expect the choice of prior copula to impact whether the posterior mode is closer to $\boldsymbol{\theta}_{0}$ for large sample sizes $n$.
@@
$\begin{split}\frac{\partial}{\partial\theta_{j}}\bigg{[}\text{log}(c_{2}(\boldsymbol{u}))-\text{log}(c_{1}(\boldsymbol{u}))\bigg{]}=\\ f_{j}(\theta_{j})\frac{\partial}{\partial u_{j}}\bigg{[}\text{log}(c_{2}(\boldsymbol{u}))-\text{log}(c_{1}(\boldsymbol{u}))\bigg{]}.\end{split}$ (8)
We can factor out the $f_{j}(\theta_{j})$ term because $p_{1}(\boldsymbol{\theta})$ and $p_{2}(\boldsymbol{\theta})$ are defined using the same marginals. The priors were also defined such that $p_{1}(\boldsymbol{\theta})>0$ and $p_{2}(\boldsymbol{\theta})>0$ for all $\boldsymbol{\theta}\in\boldsymbol{\Theta}$. Therefore, $f_{j}(\theta_{j})$ must be positive, and the partial derivative in (8) with respect to $\theta_{j}$ is 0 if and only if the partial derivative with respect to $u_{j}$ is 0. This correspondence ensures the results from Theorem 2 generalize over different specifications for the marginal distributions that are common to both priors. Counterparts to Theorem 2 for settings where both the marginals and copulas are allowed to differ between $p_{1}(\boldsymbol{\theta})$ and $p_{2}(\boldsymbol{\theta})$ would be less broadly applicable and more conceptually complex.